# Spectral Structure of the Mixed Hessian of the Dispersionless Toda $\tau$-Function

### 24 Mar 2026 v1

@@ 5.1. Intrinsic Hessian, weighted realization, and continued scalar data
Three related objects must be kept distinct beyond the analytic threshold. First, the intrinsic Hessian coefficients $H_{mn}$ and the corresponding unweighted block coefficients are attached to the conformal map itself and therefore remain meaningful as long as the geometric map exists. Second, the weighted block operators $\widetilde{G}^{(q)}(\zeta)$ from Section 4, constructed for a fixed choice of $\beta>0$, are a fixed-space Hilbert realization of the subcritical theory. Their role is to separate the logarithmically singular rank-one part from the bounded soft remainder. Third, the scalar quantities $\mathcal{G}_{p}$, $\sigma_{p}$, and $\rho_{p}$ encode the same branch-point information at the coefficient level and are the objects that can be continued beyond the weighted operator regime.
The weighted compact-operator framework of Section 4 is inherently subcritical. As $\zeta\uparrow\zeta_{c}$, the scalar Gram weights $\sigma_{p}(\zeta)$ acquire the borderline logarithmic divergence isolated in Theorem 4.6. Accordingly, the fixed-space weighted block realization used in Section 4 is not continued past $\zeta_{c}$: no bounded continuation of the compact operators $\widetilde{G}^{(q)}(\zeta)$ on the same weighted Hilbert space is asserted here.
What does continue is the scalar data built from the squared Raney coefficients. The purpose of this section is to isolate those scalar quantities, continue them analytically beyond the radius of convergence, and show that they still encode the branch-point mechanism responsible for the subcritical spectral instability. Thus this is a scalar continuation theory, not an extension of the weighted operator picture beyond $\zeta_{c}$. Since the geometric map remains univalent up to the larger threshold $\zeta_{\mathrm{univ}}$; see Proposition 6.1, the interval
$\zeta_{c}<\zeta<\zeta_{\mathrm{univ}}$
should be understood as a regime in which the fixed-space weighted spectral realization breaks down before geometric univalence is lost.
The argument has three parts. First, we introduce the scalar functions $\mathcal{G}_{p}(u)$, where $u=\zeta^{2}$, and show that they are generalized hypergeometric functions $_{2s}F_{2s-1}$ with parametric excess $\gamma=2$; see Proposition 5.4 and Corollary 5.5. This yields a canonical analytic continuation to the slit plane
$\mathbb{C}\setminus[\zeta_{c}^{2},\infty).$
Second, we analyze the singular point $u=\zeta_{c}^{2}$ and prove the resonant local expansion
$\mathcal{G}_{p}(u)=A(u)+B(u)\Bigl(1-\frac{u}{\zeta_{c}^{2}}\Bigr)^{2}\log\!\Bigl(1-\frac{u}{\zeta_{c}^{2}}\Bigr),$
@@
with $A$ and $B$ analytic and $B(\zeta_{c}^{2})<0$; see Theorem 5.9. The Euler operator that reconstructs the scalar Gram weights from $\mathcal{G}_{p}$ removes the quadratic prefactor and thereby converts this mild resonant singularity into the logarithmic divergence found in Section 4; see Corollary 5.10. Third, the branch-cut discontinuity of $\mathcal{G}_{p}$ gives a Cauchy–Stieltjes representation; see Proposition 5.16. In the range $1\leq p\leq s$, the representing measure is positive, and $\mathcal{G}_{p}$ becomes the Weyl $m$-function of a bounded Jacobi operator $J_{p}$; see Proposition 5.19 and Theorem 5.21.
@@
Throughout this section we work with the continuation variable $u=\zeta^{2}$. Later, in the Stieltjes/Jacobi interpretation, we shall also use the reciprocal spectral variable $t=u^{-1}$. To avoid confusion, we reserve $\rho_{p}(u)$ for the jump density of the continued scalar Gram weight and $\varrho_{p}(t)$ for the density of the representing measure of $\mathcal{G}_{p}$. For a function $f$ defined on the slit plane, we write
@@ 5.2. From Gram weights to scalar generating functions
We begin by separating the genuinely analytic part of the Gram weights from the polynomial prefactor coming from the block index. Recall from Definition 3.7 that
@@
The point of Definition 5.1 is that the original Gram weight is recovered from $\mathcal{G}_{p}$ by a fixed second-order Euler operator. Thus the continuation problem for $\sigma_{p}$ reduces to the continuation problem for $\mathcal{G}_{p}$.
@@ 5.3. Hypergeometric structure and analytic continuation
We next identify the analytic structure of the scalar generating function $\mathcal{G}_{p}$ from Definition 5.1. The first point is that its radius of convergence is exactly the analytic threshold $u=\zeta_{c}^{2}$, which is already encoded in the large-$m$ asymptotics of the Raney coefficients.
@@
The representation (5.2) is local in $u$. To continue $\mathcal{G}_{p}$ beyond $|u|<\zeta_{c}^{2}$, one needs a closed analytic model. This is provided by a generalized hypergeometric expression, whose existence ultimately reflects the fact that the coefficient ratio $R_{s,p}(m+1)^{2}/R_{s,p}(m)^{2}$ is rational in $m$.
@@
(5.6) $\{\alpha_{i}\}=2\times\left\{\frac{p+k}{s}:k=0,\ldots,s-1\right\},\qquad\{\beta_{j}\}=\{1\}\ \cup\ 2\times\left\{\frac{p+l}{s-1}:l=1,\ldots,s-1\right\}.$
@@
Lemma 5.6 (Reduced hypergeometric data after cancellation).
@@
If the reduction were of type $_{1}F_{0}$, its unique upper parameter would have to equal 2-2 in order to have excess $2$, which is impossible because every surviving upper parameter comes from the positive list (5.6). Thus $q_{p}\geq 1$, so the reduced form is genuinely of type $_{q_{p}+1}F_{q_{p}}$. For such equations the only finite singular points are 0 and $1$ in the variable $\xi$; see [18, §16.8, §16.11]. ∎
@@
By Lemma 5.6, after cancellation of common upper and lower parameters the germ at $u=0$ is represented by a reduced generalized hypergeometric function of type $_{q_{p}+1}F_{q_{p}}$ with the same parametric excess $2$. For that reduced hypergeometric differential equation, the only finite singular points in the variable $\xi=\zeta_{c}^{-2}u$ are $\xi=0$ and $\xi=1$; equivalently, the only finite singular points in the variable $u$ are $u=0$ and $u=\zeta_{c}^{2}$; see [18, §16.8, §16.11]. Hence the germ at the origin continues uniquely along every path in $\mathbb{C}\setminus[\zeta_{c}^{2},\infty)$, which is simply connected. This yields a single-valued holomorphic continuation of $\mathcal{G}_{p}$ to the slit plane. ∎
@@ 5.4. Resonant expansion at the branch point
We now analyze the singularity of $\mathcal{G}_{p}$ at the endpoint $u=\zeta_{c}^{2}$ of the disk of convergence. Because the parametric excess is the integer $\gamma=2$, the local Frobenius basis at the branch point is resonant, and the singular term is of the form $(1-u/\zeta_{c}^{2})^{2}\log(1-u/\zeta_{c}^{2})$.
Lemma 5.8 (Reduced local model at the branch point).
@@
$\{0,2\}\quad\text{if }q_{p}=1,\qquad 0,1,\dots,q_{p}-1,2\quad\text{if }q_{p}\geq 2.$
In particular, the exponent 0 is simple, while the exponent $1$ is either absent or simple. Hence no terms of the form $\log(1-\xi)$ or $(1-\xi)\log(1-\xi)$ can occur in the local continuation of the germ.
@@
By Lemma 5.6, after cancellation $\mathcal{G}_{p}$ satisfies a reduced generalized hypergeometric equation of type $_{q_{p}+1}F_{q_{p}}$ with parametric excess $\gamma=2$. For such an equation, the local exponents at $\xi=1$ are $0,1,\dots,q_{p}-1,\gamma$; see [18, §16.8]. Substituting $\gamma=2$ gives the stated list. The exponent 0 is therefore simple, and the exponent $1$ is either absent or simple. Consequently, the local continuation of the germ selected by the Taylor series at $\xi=0$ cannot contain logarithmic terms of order $(1-\xi)^{0}$ or $(1-\xi)^{1}$. The first order at which a logarithmic sector may occur is thus $(1-\xi)^{2}$. ∎
@@
as $u\to\zeta_{c}^{2}$ in $\mathbb{C}\setminus[\zeta_{c}^{2},\infty)$. Moreover,
@@
Set $\xi:=u/\zeta_{c}^{2}$. By Corollary 5.7, the germ defined by the Taylor series at $\xi=0$ extends holomorphically to the slit plane $\mathbb{C}\setminus[1,\infty)$. By Lemma 5.6, after cancellation of any common upper and lower parameters in (5.5), this germ is a reduced generalized hypergeometric function of type $_{q_{p}+1}F_{q_{p}}$ with parametric excess
@@
By Lemma 5.8, the local exponents at $\xi=1$ are 0 and $1$ with no multiplicity, so logarithmic terms of order $(1-\xi)^{0}$ and $(1-\xi)^{1}$ are excluded. For a reduced generalized hypergeometric equation of type $_{q_{p}+1}F_{q_{p}}$ with positive integer parametric excess $m$, the connection formulas at $\xi=1$ on the slit plane imply that the continued germ has a local representation consisting of an analytic part plus one resonant term of the form $(1-\xi)^{m}\log(1-\xi)$ with analytic coefficient; see [18, §16.8(ii), §16.11]. Since here $m=\gamma=2$, it follows that
@@
where $A_{0}$ and $B_{0}$ are analytic near $\xi=1$. The exponent analysis from Lemma 5.8 shows that no lower-order logarithmic terms occur.
Returning to the variable $u=\zeta_{c}^{2}\xi$, we obtain (5.7) with
@@
The coefficient $B(\zeta_{c}^{2})$ is computed explicitly in Appendix D. More precisely, Proposition D.1 yields
$B(\zeta_{c}^{2})<0$
for all $s\geq 2$ and $p\geq 1$. Hence the logarithmic sector is nontrivial and starts precisely at order $(1-\xi)^{2}\log(1-\xi)$. ∎
The singularity in (5.7) is mild: the function $\mathcal{G}_{p}(u)$ itself stays finite at $u=\zeta_{c}^{2}$. The logarithmic divergence appears only after one returns to the Gram weights via the Euler operator from Lemma 5.2. In this sense, the scalar continuation problem is strictly softer than the operator-theoretic one.
Corollary 5.10 (Mechanism of Gram-weight divergence).
@@ 5.5. Continued Gram weights and edge density
The Euler identity from Lemma 5.2 continues to make sense on the slit plane once $\mathcal{G}_{p}$ has been analytically continued. This gives a canonical continuation of the scalar Gram weights.
Definition 5.11 (Continued scalar Gram weights).
@@
For $|u|<\zeta_{c}^{2}$, this agrees with the original Gram weight $\sigma_{p}(\zeta)$ under the substitution $u=\zeta^{2}$, by Lemma 5.2.
@@
To illustrate the continuation across the critical point, we plot in Figure 3 the continued scalar Gram weight $\sigma_{p}^{\mathrm{cont}}(u)$, which agrees with the subcritical Gram weight for $u<\zeta_{c}^{2}$, together with the supercritical discontinuity density $\rho_{p}(u)$ for several values of $p$. The lower panels show that the edge value
@@
is positive, in agreement with the local expansion at the branch point. At the same time, the additional curves with larger $p$ show that $\rho_{p}(u)$ need not remain positive for all $u>\zeta_{c}^{2}$: after starting from a positive edge value, the density may cross zero and become negative further along the supercritical branch. Thus the sign of $\rho_{p}$ is controlled locally near $u=\zeta_{c}^{2}$, but is not globally fixed on the whole continuation domain.
Figure 3. Continued Gram weight $\sigma_{p}^{\mathrm{cont}}(u)$ and discontinuity density $\rho_{p}(u)$ for $s=3,5$. Top row: the analytically continued Gram weight $\sigma_{p}^{\mathrm{cont}}(u)$ on the subcritical side $0<u<\zeta_{c}^{2}$. Bottom row: the discontinuity density $\rho_{p}(u)$ on the supercritical side $u>\zeta_{c}^{2}$. The open circles at $u=\zeta_{c}^{2}$ mark the edge values $\rho_{p}(\zeta_{c}^{2})>0$. For larger values of $p$, the density becomes negative away from the edge, showing that positivity at $u=\zeta_{c}^{2}$ does not persist on the entire supercritical branch.
@@ 5.6. Cauchy–Stieltjes representation
each occurring with multiplicity $2$. The generalized hypergeometric differential equation has regular singular points at $\xi=0,1,\infty$, and the local exponents at $\xi=\infty$ are precisely the numerator parameters; see [18, §16.11]. Because each exponent appears with multiplicity $2$, the asymptotic expansion contains at most one logarithm per exponent:
@@
Since $\mathcal{G}_{p}$ is holomorphic on $\mathbb{C}\setminus[\zeta_{c}^{2},\infty)$, the function $F_{p}$ is holomorphic on $\mathbb{C}\setminus[0,1/\zeta_{c}^{2}]$. From the Taylor expansion $\mathcal{G}_{p}(u)=1+O(u)$ at $u=0$, one obtains
$F_{p}(\eta)=\eta^{-1}+O(\eta^{-2}),\qquad|\eta|\to\infty.$
On the other hand, Lemma 5.15 gives
$F_{p}(\eta)=o(|\eta|^{-1}),\qquad\eta\to 0,$
away from the cut.
@@
For general $p$, the representation (5.16)–(5.17) is used only as an analytic description of the continued scalar Gram data. Positivity is not required at this stage. The operator-theoretic interpretation enters only later, in the range $1\leq p\leq s$, where $\nu_{p}$ becomes a positive measure and one can pass to the Jacobi/OPRL framework.

### 16 Apr 2026 v2

@@ 5. Scalar Gram continuation beyond the analytic threshold
Beyond the analytic threshold, three layers of structure must be kept separate. The intrinsic Hessian coefficients $H_{mn}$ and the corresponding unweighted block coefficients remain attached to the conformal map itself and therefore make sense as long as the geometric map exists. By contrast, the weighted block operators $\widetilde{G}^{(q)}(\zeta)$ from Section 4 are a fixed-space Hilbert realization of the subcritical theory, introduced to isolate the logarithmically singular rank-one direction from the bounded soft remainder. The scalar quantities $\mathcal{G}_{p}$, $\sigma_{p}$, and $\rho_{p}$ reflect the same branch-point mechanism at coefficient level and are the objects that continue naturally beyond the weighted operator regime.
Accordingly, the weighted compact-operator framework is not continued past $\zeta_{c}$. As $\zeta\uparrow\zeta_{c}$, the Gram weights $\sigma_{p}(\zeta)$ already acquire the borderline logarithmic divergence isolated in Theorem 4.6. What persists beyond the radius of convergence is the scalar quantities built from the squared Raney coefficients. The purpose of this section is to continue those scalar quantities across the slit plane $\mathbb{C}\setminus[\zeta_{c}^{2},\infty)$, identify the branch-point singularity at $u=\zeta_{c}^{2}$, and recover from it the subcritical logarithmic divergence.
@@
The section proceeds in three steps. We first identify $\mathcal{G}_{p}(u)$ as a generalized hypergeometric function and obtain its canonical continuation; see Proposition 5.4 and Corollary 5.5. We then analyze the singular point $u=\zeta_{c}^{2}$ and prove the resonant expansion from Theorem 5.9. Applying the Euler operator that reconstructs the Gram weights removes the quadratic prefactor and recovers the logarithmic divergence from Section 4; see Corollary 5.10. Finally, the branch-cut jump of $\mathcal{G}_{p}$ yields a Cauchy–Stieltjes representation, and in the range $1\leq p\leq s$ this representation is positive and hence admits a Jacobi-matrix realization; see Proposition 5.16, Proposition 5.19, and Theorem 5.21.
@@
Throughout this section we use the continuation variable $u=\zeta^{2}$. In the Stieltjes/Jacobi interpretation we also use the reciprocal spectral variable $t=u^{-1}$. Accordingly, we reserve $\rho_{p}(u)$ for the jump density of the continued Gram weights and $\varrho_{p}(t)$ for the density of the representing measure of $\mathcal{G}_{p}$. For a function $f$ defined on the slit plane, we write
@@ 5.1. From Gram weights to scalar generating functions
We now isolate the scalar object that carries the singularity. Recall from Definition 3.7 that
@@
Definition 5.1 removes the explicit polynomial prefactor and leaves a scalar series whose analytic continuation can be studied directly. The original Gram weights are recovered from $\mathcal{G}_{p}$ by a fixed second-order Euler operator.
@@ 5.2. Hypergeometric structure and analytic continuation
We next identify the analytic representation for $\mathcal{G}_{p}$. First, its radius of convergence is exactly the analytic threshold $u=\zeta_{c}^{2}$, already encoded in the large-$m$ asymptotics of the Raney coefficients.
@@
To continue $\mathcal{G}_{p}$ beyond $|u|<\zeta_{c}^{2}$, it is enough to exploit that the coefficient ratio $R_{s,p}(m+1)^{2}/R_{s,p}(m)^{2}$ is rational in $m$. This places $\mathcal{G}_{p}$ in the generalized hypergeometric class.
@@
(5.6) $\{\alpha_{i}\}=2\times\left\{\frac{p+k}{s}:k=0,\ldots,s-1\right\},\quad\{\beta_{j}\}=\{1\}\ \cup\ 2\times\left\{\frac{p+l}{s-1}:l=1,\ldots,s-1\right\}.$
@@
Lemma 5.6 (Reduced hypergeometric parameters after cancellation).
@@
If the reduction were of type $_{1}F_{0}$, its unique upper parameter would have to equal 2-2 in order to have excess $2$, which is impossible because every surviving upper parameter comes from the positive list (5.6). Thus $q_{p}\geq 1$, so the reduced form is genuinely of type $_{q_{p}+1}F_{q_{p}}$. For such equations the only finite singular points are 0 and $1$ in the variable $\xi$; see [17, §16.8, §16.11]. ∎
@@
By Lemma 5.6, after cancellation of common upper and lower parameters the germ at $u=0$ is represented by a reduced generalized hypergeometric function of type $_{q_{p}+1}F_{q_{p}}$ with the same parametric excess $2$. For that reduced hypergeometric differential equation, the only finite singular points in the variable $\xi=\zeta_{c}^{-2}u$ are $\xi=0$ and $\xi=1$. Equivalently, the only finite singular points in the variable $u$ are $u=0$ and $u=\zeta_{c}^{2}$; see [17, §16.8, §16.11]. Hence the germ at the origin continues uniquely along every path in $\mathbb{C}\setminus[\zeta_{c}^{2},\infty)$, which is simply connected. This yields a single-valued holomorphic continuation of $\mathcal{G}_{p}$ to the slit plane. ∎
@@ 5.3. Resonant expansion at the branch point
We now analyze the singularity of $\mathcal{G}_{p}$ at the endpoint $u=\zeta_{c}^{2}$ of the disk of convergence. Because the parametric excess is the integer $\gamma=2$, the local continuation at the branch point is resonant, and the first possible logarithmic term occurs at order $(1-u/\zeta_{c}^{2})^{2}\log(1-u/\zeta_{c}^{2})$.
Lemma 5.8 (Reduced local description at the branch point).
@@
$\{0,2\}\quad\text{if }q_{p}=1,\qquad\{0,1,2\}\quad\text{if }q_{p}=2,\qquad\{0,1,\dots,q_{p}-1\}\cup\{2\}\quad\text{if }q_{p}\geq 3,$
where in the last case the exponent $2$ has multiplicity two. In particular, the exponent 0 is simple, while the exponent $1$ is either absent or simple. Hence no terms of the form $\log(1-\xi)$ or $(1-\xi)\log(1-\xi)$ can occur in the local continuation of the germ.
@@
By Lemma 5.6, after cancellation $\mathcal{G}_{p}$ satisfies a reduced generalized hypergeometric equation of type $_{q_{p}+1}F_{q_{p}}$ with parametric excess $\gamma=2$. For such an equation, the local exponents at $\xi=1$ are $0,1,\dots,q_{p}-1,\gamma$; see [17, §16.8]. Substituting $\gamma=2$ gives the stated list, with exponent $2$ repeated when $q_{p}\geq 3$. The exponent 0 is therefore simple, and the exponent $1$ is either absent or simple. Consequently, the local continuation of the germ selected by the Taylor series at $\xi=0$ cannot contain logarithmic terms of order $(1-\xi)^{0}$ or $(1-\xi)^{1}$. The first order at which a logarithmic term may occur is thus $(1-\xi)^{2}$. ∎
@@
for all $u$ in some slit neighborhood of $\zeta_{c}^{2}$ contained in $\mathbb{C}\setminus[\zeta_{c}^{2},\infty)$. Moreover,
@@
Set $\xi:=u/\zeta_{c}^{2}$. By Corollary 5.7, the Taylor germ at $\xi=0$ extends holomorphically to $\mathbb{C}\setminus[1,\infty)$. By Lemma 5.6, after cancellation of common upper and lower parameters in (5.5), this continuation is a reduced $_{q_{p}+1}F_{q_{p}}-$function with parametric excess
@@
Hence $\xi=1$ is a regular singular point of the corresponding differential equation. By Lemma 5.8, the exponents 0 and $1$ at $\xi=1$ are simple, so no logarithmic terms of order $(1-\xi)^{0}$ or $(1-\xi)^{1}$ can occur in the continuation of the Taylor germ from $\xi=0$. The remaining Frobenius sectors, when present, start at higher integer powers and are analytic at $\xi=1$. For a reduced $_{q_{p}+1}F_{q_{p}}-$equation with positive integer parametric excess $m$, the analytic continuation of the Taylor germ at $\xi=0$ has a local representation near $\xi=1$ of the form
$(1-\xi)^{m}\bigl(C_{0}(\xi)+D_{0}(\xi)\log(1-\xi)\bigr),$
where $C_{0}$ and $D_{0}$ are analytic near $\xi=1$; see [17, §16.8(i), §16.11(ii)] and [22, Ch. 4]. Since here $m=\gamma=2$, we obtain on a slit neighborhood of $\xi=1$
@@
where $A_{0}$ and $B_{0}$ are analytic near $\xi=1$. The simplicity of the exponents 0 and $1$ excludes lower-order logarithmic terms. Returning to $u=\zeta_{c}^{2}\xi$, define
@@
This gives (5.7). The sign of $B(\zeta_{c}^{2})$ is computed independently in Proposition D.1, which gives
$B(\zeta_{c}^{2})=-\frac{p^{2}}{4\pi}\,\frac{s^{2p-1}}{(s-1)^{2p+1}}<0.$
Corollary 5.10 (Mechanism of Gram weights divergence).
@@ 5.4. Continued Gram weights and edge density
Once $\mathcal{G}_{p}$ has been continued to the slit plane, the Euler identity from Lemma 5.2 gives a canonical continuation of the Gram weights as well.
Definition 5.11 (Continued Gram weights).
@@
For $|u|<\zeta_{c}^{2}$, this agrees with the original Gram weights $\sigma_{p}(\zeta)$ under the substitution $u=\zeta^{2}$, by Lemma 5.2.
@@
To illustrate the continuation across the critical point, we plot in Figure 3 the continued Gram weights $\sigma_{p}^{\mathrm{cont}}(u)$, which agrees with the subcritical Gram weights for $u<\zeta_{c}^{2}$, together with the supercritical discontinuity density $\rho_{p}(u)$ for several values of $p$. The lower plots show that the edge value
@@
is positive, in agreement with the local expansion at the branch point. At the same time, the additional curves with larger $p$ show that $\rho_{p}(u)$ need not remain positive for all $u>\zeta_{c}^{2}$: after starting from a positive edge value, the density may cross zero and become negative further along the supercritical branch. Thus the sign of $\rho_{p}$ is determined locally near $u=\zeta_{c}^{2}$, but need not remain fixed throughout the whole continuation domain.
Figure 3. Continued Gram weights $\sigma_{p}^{\mathrm{cont}}(u)$ and discontinuity density $\rho_{p}(u)$ for $s=3,5$. Top row: the analytically continued Gram weights $\sigma_{p}^{\mathrm{cont}}(u)$ on the subcritical side $0<u<\zeta_{c}^{2}$. Bottom row: the discontinuity density $\rho_{p}(u)$ on the supercritical side $u>\zeta_{c}^{2}$. The open circles at $u=\zeta_{c}^{2}$ mark the edge values $\rho_{p}(\zeta_{c}^{2})>0$. For larger values of $p$, the density becomes negative away from the edge, showing that positivity at $u=\zeta_{c}^{2}$ does not persist on the entire supercritical branch.
@@ 5.5. Cauchy–Stieltjes representation
each occurring with multiplicity $2$. The generalized hypergeometric differential equation has regular singular points at $\xi=0,1,\infty$, and the local exponents at $\xi=\infty$ are precisely the numerator parameters; see [17, §16.11(i)] and [22, Ch. 4]. Because each exponent appears with multiplicity $2$, the local basis at $\infty$ contains at most one logarithm for each exponent, so the asymptotic expansion takes the form
@@
Since $\mathcal{G}_{p}$ is holomorphic on $\mathbb{C}\setminus[\zeta_{c}^{2},\infty)$, the function $F_{p}$ is holomorphic on $\mathbb{C}\setminus[0,1/\zeta_{c}^{2}]$. From the Taylor expansion $\mathcal{G}_{p}(u)=1+O(u)$ at $u=0$, one obtains $F_{p}(\eta)=\eta^{-1}+O(\eta^{-2})$, as $|\eta|\to\infty$. On the other hand, Lemma 5.15 gives $F_{p}(\eta)=o(|\eta|^{-1})$, as $\eta\to 0$, away from the cut.
@@
For general $p$, the representation (5.16)–(5.17) is used only as an analytic description of the continued scalar Gram quantities. Positivity is not required at this stage. The operator-theoretic interpretation enters only later, in the range $1\leq p\leq s$, where $\nu_{p}$ becomes a positive measure and one can pass to the Jacobi/OPRL framework.